jigglypuff314 (jigglypuff314):

I need help with a Differential Equations problem. Chapter on undamped forcing and resonance...

jigglypuff314 (jigglypuff314):

\[\frac{ d^2y }{ dt^2 } + 9y = 2 \cos(3t)\]

jigglypuff314 (jigglypuff314):

Compute the general solution of the given equation

jigglypuff314 (jigglypuff314):

I'm not sure if notations hold true in general but I have an idea that \(\large y(t) = y_h(t)+y_p(t)\)

jigglypuff314 (jigglypuff314):

I think I got that ... \(y_h(t) = k_1 cos(3t) + k_2 sin(3t)\)

jigglypuff314 (jigglypuff314):

anyways, I am stuck on getting \(y_p(t)\) using complexification or otherwise .-.

OpenStudy (just_one_last_goodbye):

Give me a sec ^.^

Elsa213 (elsa213):

Use a calculator e.e

jigglypuff314 (jigglypuff314):

this is a practice problem for the quiz I'm gonna have to take later today - no calculators allowed. I need to know the correct steps and processes and logic so I can do well on the quiz.

Elsa213 (elsa213):

@.Sam. @IrishBoy123 @iYuko

jigglypuff314 (jigglypuff314):

@Directrix @skullpatrol I require math help

OpenStudy (just_one_last_goodbye):

I got the answer Jiggly :D

Elsa213 (elsa213):

@.Sam. the smart one needs the smartest one. e.e

OpenStudy (just_one_last_goodbye):

Ok Ok... but it's long .-.

ILovePuppiesLol (ilovepuppieslol):

@inkyvoyd

OpenStudy (just_one_last_goodbye):

Wait you only want the yp(t) right?

jigglypuff314 (jigglypuff314):

yes

ILovePuppiesLol (ilovepuppieslol):

stop spamming, i doubt you can do this, if you are not going to help, leave @just_one_last_goodbye

jigglypuff314 (jigglypuff314):

^

jigglypuff314 (jigglypuff314):

I have so far attempted \(\large y_p(t) = ke^{3it}\) and \(\large y_p(t) = a*cos(3t) + b*sin(3t)\) both of which led to dead ends

OpenStudy (just_one_last_goodbye):

Heres what you want Jiggly >.> http://prntscr.com/dh15cm

OpenStudy (just_one_last_goodbye):

Let me know if my calculator is failing your needs .-. so I may improve it

jigglypuff314 (jigglypuff314):

I did that and everything on the left side seemed to cancel out

OpenStudy (518nad):

y_p(t), means a particular solution

jigglypuff314 (jigglypuff314):

yes

OpenStudy (518nad):

take y(t)=Acos(kt)

OpenStudy (just_one_last_goodbye):

http://prntscr.com/dh16zg

ILovePuppiesLol (ilovepuppieslol):

leave @just_one_last_goodbye

jigglypuff314 (jigglypuff314):

I see one of my mistakes now... brb I'ma recalculate

OpenStudy (just_one_last_goodbye):

@ILovePuppiesLol leave me alone bruh ;-; im trying to help... not spam like you

jigglypuff314 (jigglypuff314):

I forgot a chain rule...

OpenStudy (just_one_last_goodbye):

General solution.... \[y = c1\cos (3t) + c2\sin (3t) + \frac{ 1 }{ 3 } tsin (3t)\]

OpenStudy (518nad):

are you good now?

jigglypuff314 (jigglypuff314):

I told you know already knew the answer

jigglypuff314 (jigglypuff314):

I needed to know HOW to get it so I'm still trying to work that out after finding my mistake, brb

OpenStudy (just_one_last_goodbye):

._. my calculator has helped :D

OpenStudy (just_one_last_goodbye):

Sadly Preetha rejected it :( after all my hard work trying to make it a part of OS...

jigglypuff314 (jigglypuff314):

I'm stuck at -8at + 6b = 2 -6a - 8bt = 0

OpenStudy (518nad):

why is there a t

jigglypuff314 (jigglypuff314):

when using \(\large y_p(t) = at*cos(3t) + bt*sic(3t)\)

jigglypuff314 (jigglypuff314):

*sin(3t)

OpenStudy (518nad):

dont use undetermined coeff the constants are cancelling out, they did it on purpose, u cant do it without some trig identities

OpenStudy (518nad):

use the general variation of parameter form

OpenStudy (518nad):

once you find your y_h you can find your y_p with this formula http://prntscr.com/dh1akg

jigglypuff314 (jigglypuff314):

interesting... I've never learned that

OpenStudy (518nad):

so you've only learn undetermined coefficients?

OpenStudy (518nad):

u can still do it that way, you just need to do a bit of algebra

jigglypuff314 (jigglypuff314):

and complexification

OpenStudy (518nad):

okok so then do this

jigglypuff314 (jigglypuff314):

\(\large y_p(t) = ke^{i \omega t} \)

jigglypuff314 (jigglypuff314):

I figured out how -8at + 6b = 2 -6a - 8bt = 0 gets me the right answer .-. not sure why I had to set t to 0 but ok

OpenStudy (518nad):